OpenStudy (anonymous):

What is the slope of a line that passes through a pair of points, (-5/3, -1) and (-2, 9/2) A. 2/33 B. -2/33 C. -33/2 D. 33/2

OpenStudy (anonymous):

\[\frac{ (-1)-(9/2) }{ (-5/3)-(-2) }\] \[\frac{ (-1)-4.5 }{ (-5/3)+2 }\] \[\frac{ -5.5 }{ (-5/3)+(6/3) }\] \[\frac{ -5.5}{ (1/3) }\] -16.5 now put that into a fraction